What does data distribution mean?
It is the set of values that a variable can take together with the probability or frequency distribution for those values.
What is the probability of getting a heart out of a pack of cards?
If the pick is completely random, the deck is a standard deck and there are no jokers or any other cards other than the standard 52, the probability is 1/4
Find the standard deviation of the binomial distribution for which n 1000 and p 0.94?
The standard deviation is sqrt[n*p*(1-p)] = sqrt(1000*0.94*0.06) = 7.51 approx.
What is the probability of an observation being above the mean?
For a continuous variable it is 0.5 but for a discrete variable the answer depends on the probability of the variable taking the mean value. It is half of the rest of the probability.
If the discrete variable X has mean m and Prob(X = m) is p
then Prob(X > m) = (1 - p)/2.
What is the probabilty of drawing a face card that has a diamond?
3/52
Assuming you're randomly drawing from a usual deck 52 unique cards. Jack Queen King are the face cards.
How many distinct arrangements can be made with the letters in the word banana?
There are 6!/(3!*2!) = 60 arrangements.
There are 11880 ways.
It is: 12 out of 50 or 6 out of 25
If there is a correlation between event a and Event B what statements is true?
event B has something in common with event A
Answer: 1/10 for no tickets, 6/10 for two tickets, 3/10 tickets for one ticket.
What is the probability of picking a month that starts with J given that it ends in y?
The months that end in y are January, February, May, and July. Two of those four start with J. So 2/4=.5 is the answer.
Is it true if two events are indenpent then the probability of both events is less than 1?
Not necessarily.
What is the probability that the sum of the dice is odd and both dice show the number 3?
The probability is 0. If both dice show the number 3 then the sum is 6 which is not odd.
The total probability of something happening plus the probability of that same thing not happening is 1, or 100 %
→ probability of not happening = 1 - 0.25 = 0.75 or 100 % - 25 % = 75 %
What is the probability of choosing a prime number?
Although there are infinitely many primes, they become rarer and rarer so that as the number of numbers increases, the probability that picking one of them at random is a prime number tends to zero*.
In the first 10 numbers there are 4 primes, so the probability of picking one is 4/10 = 2/5 = 0.4
In the first 100 numbers there are 26 primes, so the probability of picking one is 25/100 = 1/4 = 0.25
In the first 1,000 numbers there are 169 primes, so the probability of picking one is 168/1000 = 0.168
In the first 10,000 numbers there are 1,229 primes, so the probability of picking one is 0.1229
In the first 100,000 numbers there are 9592 primes, so the probability of picking one is 0.09592
In the first 1,000,000 numbers there are 78,498 primes, so the probability of picking one is 0.078498
In the first 10,000,000 numbers there are 664,579 primes, so the probability of picking one is 0.0664579
* Given any small value ε less than 1 and greater than 0, it is possible to find a number n such that the probability of picking a prime at random from the numbers 1-n is less than the given small value ε.
What is the probability that both cards are spades?
The answer will depend on:
It also depends on how many times the experiment - of drawing two cards - is repeated. If repeated a sufficient number of times the probability will be so close to 1 as to make no difference from a certainty.
It is 0.25
How many ways can 9 books be arranged on a shelf if there is only room for 5 of the books?
9 x 8 x 7 x 6 x 5
It is 84.3%
What is the probability to take a Jack in 52 cards?
1:13
or 4:52
as four suits and one jack in each suit
You can expect the spinner to land an odd number 25 times out of 50.